OpenStudy (anonymous):

2. Identify a pattern and find the next number in the pattern. –48, –24, –12, –6 (1 point) –3 0 2 3

OpenStudy (anonymous):

here you see -48/2=-24,-24/2=-12,-12/2=-6 so the next step will be -6/2=-3

OpenStudy (anonymous):

oh so it is by dividing so the answer is -3

OpenStudy (anonymous):

Technicallly, it isn't dividing. This is a geometric progression in which each successive term is one half of the previous term. But since multiplying by one half is the same thing as dividing by 2, you basically have your answer either way.

OpenStudy (anonymous):

thank you very much

OpenStudy (anonymous):

thank you very much